# Improved Regret Bounds for Gaussian Process Upper Confidence Bound in Bayesian Optimization

### v2

@@ Abstract
This paper addresses the Bayesian optimization problem (also referred to as the Bayesian setting of the Gaussian process bandit), where the learner seeks to minimize the regret under a function drawn from a known Gaussian process (GP). Under a Matérn kernel with a certain degree of smoothness, we show that the Gaussian process upper confidence bound (GP-UCB) algorithm achieves $\widetilde{O}(\sqrt{T})$ cumulative regret with high probability. Furthermore, our analysis yields $O(\sqrt{T\ln^{2}T})$ regret under a squared exponential kernel. These results fill the gap between the existing regret upper bound for GP-UCB and the best-known bound provided by Scarlett [43]. The key idea in our proof is to capture the concentration behavior of the input sequence realized by GP-UCB, enabling a more refined analysis of the GP’s information gain.
@@ 1 Introduction
We study the Bayesian optimization (BO) problem, where the learner seeks to minimize the regret under a random function drawn from a known Gaussian process (GP) [17, 18]. Throughout this paper, we focus on the GP-UCB algorithm [48], which combines the posterior distribution of GP with the optimism principle. Due to its simple algorithm construction and general theoretical framework provided by Srinivas et al. [48], GP-UCB has played an important role in the advancement of the BO field. On the other hand, our theoretical understanding of the performance of GP-UCB has not been improved from [48] in the Bayesian setting, while its frequentist counterpart is studied in several existing works [11, 57]. Specifically, the current regret upper bound for GP-UCB, as provided by Srinivas et al. [48], is known to be worse than that of the algorithm in [43], which achieves state-of-the-art $O(\sqrt{T\ln T})$ cumulative regret. Then, the natural question is whether there is further room for improvement in the existing regret upper bound of GP-UCB. This paper provides an affirmative answer to this question by showing that GP-UCB achieves $\widetilde{O}(\sqrt{T})$ regret with high probability.
@@
  • We show that the GP-UCB proposed by Srinivas et al. [48] achieves $\widetilde{O}(\sqrt{T})$ regret with high probability under a Matérn kernel with a certain degree of smoothness (precise condition is provided in Theorem 3). Here, $\widetilde{O}(\cdot)$ is the order notation that hides polylogarithmic dependence. This result is comparable to state-of-the-art $O(\sqrt{T\ln T})$ regret provided by Scarlett [43] up to a polylogarithmic factor and strictly improves upon the existing $\widetilde{O}(T^{\frac{\nu+d}{2\nu+d}})$ upper bound of GP-UCB [48, 54]. Here, $d$ and $\nu$ denote the dimension of the input domain and smoothness parameter, respectively.
  • Furthermore, for a squared exponential kernel, we establish $O\left(\sqrt{T\ln^{2}T}\right)$ cumulative regret of GP-UCB. This improves the existing $O\left(\sqrt{T\ln^{d+2}T}\right)$ upper bound provided by Srinivas et al. [48] for any $d\geq 1$.
@@ 1.1 Related Works
BO has been extensively studied in the past few decades. Some of them are constructed so as to maximize the utility-based acquisition function defined through the GP posterior, including expected improvement [34], knowledge gradient [16], and the entropy-based algorithms [22]. The theoretical aspect of BO has also been actively studied through the lens of the bandit algorithms, such as GP-UCB [48], Thompson sampling [40], and information directed sampling [41]. In contrast to the noisy observation setting, which these algorithms focus on, algorithms for the noise-free setting form a separate line of research [13, 21, 30]. Extensions of these algorithms to more advanced settings have also been well-studied, e.g., contextual [31], parallel observation [14], high-dimensional [27], time-varying [6], and multi-fidelity setting [28]. Unlike the Bayesian assumption on the objective function adopted in this paper, existing works also extensively study the frequentist assumption of the function, which is also referred to as the frequentist setting of BO or GP bandits [7, 9, 11, 24, 32, 42, 44, 52, 55].
Among the existing studies, [43] is closely related to this paper, which propose a successive elimination-based algorithm and shows an $O(\sqrt{T\ln T})$ upper bound and an $\Omega(\sqrt{T})$ lower bound of the cumulative regret for a one-dimensional BO problem. The fundamental theoretical assumptions and the high-level idea of our analysis are built on the proof provided by Scarlett [43]. Following [43], Wang et al. [56] also proves similar regret guarantees under the one-dimensional Brownian motion.
In addition to [43], some parts of our analysis are inspired by the technique leveraged in [8, 26]. Firstly, Cai et al. [8] studies the GP-UCB algorithm through a relaxed version of regret, which is called lenient regret. In our analysis, the cumulative regret is decomposed into the lenient regret-based term, and we leverage their technique to analyze it. Secondly, Janz et al. [26] proposed the input partitioning-based algorithm for obtaining a superior regret in the frequentist setting. Roughly speaking, the high-level idea of their analysis is based on the fact that tighter information gain bounds can be obtained within a properly shrinking partition of the input. The key idea provided in Section 3.1 is motivated by this fact, while our analysis itself is substantially different from that in [26].
@@ 2 Preliminaries
These are standard sets of assumptions in the existing theory of BO [40, 48]. Specifically, in Assumption 1, we focus on the following squared exponential (SE) kernel $k_{\mathrm{SE}}$ and Matérn kernel $k_{\mathrm{Mat\acute{e}rn}}$:
@@
Under the total step size $T\in\mathbb{N}_{+}$, the learner’s goal is to minimize the cumulative regret $R_{T}\coloneqq\sum_{t\in[T]}f(\bm{x}^{\ast})-f(\bm{x}_{t})$, where $\bm{x}^{\ast}\in\operatornamewithlimits{argmax}_{\bm{x}\in\mathcal{X}}f(\bm{x})$ and $[T]=\{1,\ldots,T\}$.
@@
The quantity $\gamma_{T}(\mathcal{X})$ is referred to as the maximum information gain (MIG) over $\mathcal{X}$ [48], since $I(\mathbf{X}_{T})$ equals the mutual information between the function values $(f(\bm{x}_{t}))_{t\in[T]}$ and the outputs $(y_{t})_{t\in[T]}$ under Assumptions 1 and 2, and the input sequence $\mathbf{X}_{t}=(\bm{x}_{1},\ldots,\bm{x}_{t})$. MIG plays a vital role in the theoretical analysis of BO, and its increasing speed is analyzed in several commonly used kernels. For example, $\gamma_{T}(\mathcal{X})=O(\ln^{d+1}T)$ as $T\rightarrow\infty$ under $k=k_{\mathrm{SE}}$ [48]. For the notational convenience, we also define $\gamma_{i}(\mathcal{X})=\gamma_{\lceil i\rceil}(\mathcal{X})$ for any non-integer $i>0$.
@@
Lemma 1 (Lipchitz condition of sample path, e.g., [48]).
Suppose $k=k_{\mathrm{SE}}$ or $k=k_{\mathrm{Mat\acute{e}rn}}$ with $\nu>2$. Assume Assumption 1. Then, there exists the constants $a,b>0$ such that
@@
Lemma 2 (Sample path condition for the global maximizer, e.g., [13, 43]).
Suppose $k=k_{\mathrm{SE}}$ or $k=k_{\mathrm{Mat\acute{e}rn}}$ with $\nu>2$. Assume Assumption 1. Then, for any $\delta_{\mathrm{GP}}\in(0,1)$, there exist the strictly positive constants $c_{\mathrm{gap}},c_{\mathrm{sup}},c_{\mathrm{lin}},c_{\mathrm{quad}},\rho_{\mathrm{lin}},\rho_{\mathrm{quad}}>0$ such that the following statements simultaneously hold with probability at least $1-\delta_{\mathrm{GP}}$:
@@
    If the maximizer $\bm{x}^{\ast}$ is at the boundary of $\mathcal{X}$, the function $f$ satisfies $\forall\bm{x}\in\mathcal{B}_{2}(\rho_{\mathrm{lin}};\bm{x}^{\ast}),~{}f(\bm{x}^{\ast})-c_{\mathrm{lin}}\|\bm{x}^{\ast}-\bm{x}\|_{2}\geq f(\bm{x})$, where $\mathcal{B}_{2}(\rho_{\mathrm{lin}};\bm{x}^{\ast})\coloneqq\{\bm{x}\in\mathcal{X}\mid\|\bm{x}^{\ast}-\bm{x}\|_{2}\leq\rho_{\mathrm{lin}}\}$ is the L2-ball on $\mathcal{X}$, whose radius and center are $\rho_{\mathrm{lin}}$ and $\bm{x}^{\ast}$, respectively.
  4. 4.
    If the maximizer $\bm{x}^{\ast}$ is a internal point of $\mathcal{X}$, the function $f$ satisfies $\forall\bm{x}\in\mathcal{B}_{2}(\rho_{\mathrm{quad}};\bm{x}^{\ast}),~{}f(\bm{x}^{\ast})-c_{\mathrm{quad}}\|\bm{x}^{\ast}-\bm{x}\|_{2}^{2}\geq f(\bm{x})$.
Lemma 1 states that the sample path $f$ of GP is a Lipschitz function with high probability. As summarized in the next paragraph, this property is leveraged in the theory of GP-UCB to control the discretization error arising from the confidence bound construction in the continuous input domain. Lemma 2 specifies the regularity condition of $f$ related to the maximizer $\bm{x}^{\ast}$. See Assumptions 2 and 3, and the discussion provided by Scarlett [43] for further details111 Compared to Assumptions 2 and 3 in [43], we omit several unnecessary conditions for our analysis. . Note that Lemma 2 is not used in the existing proof of GP-UCB in [48]. As described in the next section, we analyze the realized input sequence $\mathbf{X}_{T}$ of GP-UCB by relating it to conditions in Lemma 2.
@@
We briefly summarize the existing analysis of GP-UCB (Algorithm 1) provided by Srinivas et al. [48]. Based on Assumptions 1 and 2, we can construct the high-probability confidence bound of the underlying function value $f(\bm{x})$ for each $\bm{x}$ and $t\in\mathbb{N}_{+}$ through the posterior distribution of $f(\bm{x})$. Specifically, by choosing a properly designed finite representative input set $\mathcal{X}_{t}\subset\mathcal{X}$ and taking into account the discretization error with Lemma 1, Srinivas et al. [48] showed the following events hold simultaneously with probability at least $1-\delta$:
@@
    Discretization error. The discretization error arising from $\mathcal{X}_{t}$ is at most $1/t^{2}$. Namely, $|f(\bm{x})-f([\bm{x}]_{t})|\leq 1/t^{2}$ holds for any $\bm{x}\in\mathcal{X}$ and $t\in\mathbb{N}_{+}$, where $[\bm{x}]_{t}$ denotes one of the closest points of $\bm{x}$ on $\mathcal{X}_{t}$.
@@
In the above expression, the upper bound $\sum_{t=1}^{T}f(\bm{x}^{\ast})-f([\bm{x}^{\ast}]_{t})\leq\sum_{t=1}^{T}1/t^{2}\leq\pi^{2}/6$ follows from the second event (discretization error). The inequality $\sum_{t=1}^{T}f([\bm{x}^{\ast}]_{t})-f(\bm{x}_{t})\leq 2\beta_{T}^{1/2}\sum_{t=1}^{T}\sigma(\bm{x}_{t};\mathbf{X}_{t-1})$ also follows from the first event (confidence bound) and the definition of $\bm{x}_{t}$. See the proof of Theorem 2 in [48] for details. The above inequality suggests that the regret upper bound of GP-UCB depends on the sum of the posterior standard deviations $\sum_{t=1}^{T}\sigma(\bm{x}_{t};\mathbf{X}_{t-1})$. Srinivas et al. [48] provides the upper bound of this term by leveraging the information gain $I(\mathbf{X}_{T})$ as follows:
@@
where $C=\frac{2}{\ln(1+\sigma^{-2})}$. From Eqs. (7) and (8), we conclude that the regret upper bound of GP-UCB is $O\left(\sqrt{\beta_{T}T\gamma_{T}(\mathcal{X})}\right)$ with probability at least $1-\delta$. By combining the explicit upper bound of $\gamma_{T}(\mathcal{X})$ [48, 54], we also obtain $O\left(\sqrt{T\ln^{d+2}T}\right)$ and $\widetilde{O}\left(T^{\frac{\nu+d}{2\nu+d}}\right)$ regret upper bounds for SE and Matérn kernels, respectively.

### v3

@@ Abstract
This paper addresses the Bayesian optimization problem (also referred to as the Bayesian setting of the Gaussian process bandit), where the learner seeks to minimize the regret under a function drawn from a known Gaussian process (GP). Under a Matérn kernel with a certain degree of smoothness, we show that the Gaussian process upper confidence bound (GP-UCB) algorithm achieves $\widetilde{O}(\sqrt{T})$ cumulative regret with high probability. Furthermore, our analysis yields $O(\sqrt{T\ln^{2}T})$ regret under a squared exponential kernel. These results fill the gap between the existing regret upper bound for GP-UCB and the best-known bound provided by scarlett2018tight. The key idea in our proof is to capture the concentration behavior of the input sequence realized by GP-UCB, enabling a more refined analysis of the GP’s information gain.
@@ 1 Introduction
We study the Bayesian optimization (BO) problem, where the learner seeks to minimize the regret under a random function drawn from a known Gaussian process (GP) [frazier2018tutorial, garnett2023bayesian]. Throughout this paper, we focus on the GP-UCB algorithm [srinivas10gaussian], which combines the posterior distribution of GP with the optimism principle. Due to its simple algorithm construction and general theoretical framework provided by srinivas10gaussian, GP-UCB has played an important role in the advancement of the BO field. On the other hand, our theoretical understanding of the performance of GP-UCB has not been improved from [srinivas10gaussian] in the Bayesian setting, while its frequentist counterpart is studied in several existing works [chowdhury2017kernelized, whitehouse2023improved]. Specifically, the current regret upper bound for GP-UCB, as provided by srinivas10gaussian, is known to be worse than that of the algorithm in [scarlett2018tight], which achieves state-of-the-art $O(\sqrt{T\ln T})$ cumulative regret. Then, the natural question is whether there is further room for improvement in the existing regret upper bound of GP-UCB. This paper provides an affirmative answer to this question by showing that GP-UCB achieves $\widetilde{O}(\sqrt{T})$ regret with high probability.
@@
[…]
  • Furthermore, for a squared exponential kernel, we establish $O\left(\sqrt{T\ln^{2}T}\right)$ cumulative regret of GP-UCB. This improves the existing $O\left(\sqrt{T\ln^{d+2}T}\right)$ upper bound provided by srinivas10gaussian for any $d\geq 1$.
@@ 1.1 Related Works
BO has been extensively studied in the past few decades. Some of them are constructed so as to maximize the utility-based acquisition function defined through the GP posterior, including expected improvement [movckus1975bayesian], knowledge gradient [frazier2009knowledge], and the entropy-based algorithms [hennig2012entropy]. The theoretical aspect of BO has also been actively studied through the lens of the bandit algorithms, such as GP-UCB [srinivas10gaussian], Thompson sampling [Russo2014-learning], and information directed sampling [russo2014learning]. In contrast to the noisy observation setting, which these algorithms focus on, algorithms for the noise-free setting form a separate line of research [freitas2012exponential, grunewalder2010regret, kawaguchi2015bayesian]. Extensions of these algorithms to more advanced settings have also been well-studied, e.g., contextual [krause2011contextual], parallel observation [desautels2014parallelizing], high-dimensional [kandasamy2015high], time-varying [bogunovic2016time], and multi-fidelity setting [kandasamy2019multi]. Unlike the Bayesian assumption on the objective function adopted in this paper, existing works also extensively study the frequentist assumption of the function, which is also referred to as the frequentist setting of BO or GP bandits [bull2011convergence, camilleri2021high, chowdhury2017kernelized, iwazaki2025improvedregretanalysisgaussian, li2022gaussian, salgiarandom, scarlett2017lower, vakili2021optimal, valko2013finite].
Among the existing studies, [scarlett2018tight] is closely related to this paper, which propose a successive elimination-based algorithm and shows an $O(\sqrt{T\ln T})$ upper bound and an $\Omega(\sqrt{T})$ lower bound of the cumulative regret for a one-dimensional BO problem. The fundamental theoretical assumptions and the high-level idea of our analysis are built on the proof provided by scarlett2018tight. Following [scarlett2018tight], wang2022tight also proves similar regret guarantees under the one-dimensional Brownian motion.
In addition to [scarlett2018tight], some parts of our analysis are inspired by the technique leveraged in [cai2021lenient, janz2020bandit]. Firstly, cai2021lenient studies the GP-UCB algorithm through a relaxed version of regret, which is called lenient regret. In our analysis, the cumulative regret is decomposed into the lenient regret-based term, and we leverage their technique to analyze it. Secondly, janz2020bandit proposed the input partitioning-based algorithm for obtaining a superior regret in the frequentist setting. Roughly speaking, the high-level idea of their analysis is based on the fact that tighter information gain bounds can be obtained within a properly shrinking partition of the input. The key idea provided in Section 3.1 is motivated by this fact, while our analysis itself is substantially different from that in [janz2020bandit].
@@ 2 Preliminaries
These are standard sets of assumptions in the existing theory of BO [Russo2014-learning, srinivas10gaussian]. Specifically, in Assumption 1, we focus on the following squared exponential (SE) kernel $k_{\mathrm{SE}}$ and Matérn kernel $k_{\mathrm{Mat\acute{e}rn}}$:
@@
Under the total step size $T\in\mathbb{N}_{+}$, the learner’s goal is to minimize the cumulative regret $R_{T}\coloneqq\sum_{t\in[T]}f(\bm{x}^{\ast})-f(\bm{x}_{t})$, where $\bm{x}^{\ast}\in\operatornamewithlimits{argmax}_{\bm{x}\in\mathcal{X}}f(\bm{x})$ and $T]=\{1,\ldots,T\}$.
@@
The quantity $\gamma_{T}(\mathcal{X})$ is referred to as the maximum information gain (MIG) over $\mathcal{X}$ [srinivas10gaussian], since $I(\mathbf{X}_{T})$ equals the mutual information between the function values $(f(\bm{x}_{t}))_{t\in[T]}$ and the outputs $(y_{t})_{t\in[T]}$ under Assumptions 1 and 2, and the input sequence $\mathbf{X}_{t}=(\bm{x}_{1},\ldots,\bm{x}_{t})$. MIG plays a vital role in the theoretical analysis of BO, and its increasing speed is analyzed in several commonly used kernels. For example, $\gamma_{T}(\mathcal{X})=O(\ln^{d+1}T)$ as $T\rightarrow\infty$ under $k=k_{\mathrm{SE}}$ [srinivas10gaussian]. For the notational convenience, we also define $\gamma_{i}(\mathcal{X})=\gamma_{\lceil i\rceil}(\mathcal{X})$ for any non-integer $i>0$.
@@
Lemma 1 (Lipchitz condition of sample path, e.g., [srinivas10gaussian]).
Suppose $k=k_{\mathrm{SE}}$ or $k=k_{\mathrm{Mat\acute{e}rn}}$ with $\nu>2$. Assume Assumption 1. Then, there exist the constants $a,b>0$ such that
@@
Lemma 2 (Sample path condition for the global maximizer, e.g., [de2012regret, freitas2012exponential, scarlett2018tight]).
Suppose $k=k_{\mathrm{SE}}$ or $k=k_{\mathrm{Mat\acute{e}rn}}$ with $\nu>2$. Assume Assumption 1. Then, for any $\delta_{\mathrm{GP}}\in(0,1)$, there exist the strictly positive constants $c_{\mathrm{gap}},c_{\mathrm{sup}},c_{\mathrm{quad}},\rho_{\mathrm{quad}}>0$ such that the following statements simultaneously hold with probability at least $1-\delta_{\mathrm{GP}}$:
@@
    The function $f$ satisfies $\forall\bm{x}\in\mathcal{B}_{2}(\rho_{\mathrm{quad}};\bm{x}^{\ast}),~f(\bm{x}^{\ast})-c_{\mathrm{quad}}\|\bm{x}^{\ast}-\bm{x}\|_{2}^{2}\geq f(\bm{x})$, where $\mathcal{B}_{2}(\rho_{\mathrm{quad}};\bm{x}^{\ast})\coloneqq\{\bm{x}\in\mathcal{X}\mid\|\bm{x}^{\ast}-\bm{x}\|_{2}\leq\rho_{\mathrm{quad}}\}$ is the L2-ball on $\mathcal{X}$, whose radius and center are $\rho_{\mathrm{quad}}$ and $\bm{x}^{\ast}$, respectively [^1] In the earlier version and conference version of the paper, we claim that the linear condition holds when $\bm{x}^{\ast}$ is on the boundary of $\mathcal{X}$. However, the linear condition as in [scarlett2018tight] is only valid under $d=1$, since the partial derivative parallel to the boundary may become zero unless $\bm{x}^{\ast}$ is on the corner of $\mathcal{X}$. Therefore, for $d\geq 2$, even when $\bm{x}^{\ast}$ is on the boundary, we also have to consider the quadratic condition as with the proof of Theorem 5 in [de2012regret]. This modification only affects the constant factors in our regrets.
Lemma 1 states that the sample path $f$ of GP is a Lipschitz function with high probability. This property is leveraged in the theory of GP-UCB to control the discretization error arising from the confidence bound construction in the continuous input domain. As described in [srinivas10gaussian], Lemma 1 is a direct consequence of Theorem 5 in [ghosal2006posterior] under the existence of fourth-order mixed partial derivatives of the kernel, which are satisfied under $k=k_{\mathrm{SE}}$ and $k=k_{\mathrm{Mat\acute{e}rn}}$ with $\nu>2$ [^2] Differentiability of $k_{\mathrm{Mat\acute{e}rn}}$ is derived in the existing works, e.g., Chapter 2.7 in [stein1999interpolation].. Lemma 2 specifies the regularity condition of $f$ related to the maximizer $\bm{x}^{\ast}$. Here, property 1 is implied from the fact that the GP-sample path has a unique maximizer almost surely under $k_{\mathrm{SE}}$ and $k_{\mathrm{Mat\acute{e}rn}}$ [e.g., Lemma 2.6 in kim1990cube]. Property 2 is implied from, e.g., the compactness of $\mathcal{X}$ and the almost-sure continuity of the sample path under $k_{\mathrm{SE}}$ and $k_{\mathrm{Mat\acute{e}rn}}$. Property 3 also holds automatically under $k=k_{\mathrm{SE}}$ and $k=k_{\mathrm{Mat\acute{e}rn}}$ with $\nu>2$ and is used in existing works. See Theorem 5 in [de2012regret], Assumption 3 in [scarlett2018tight], and the discussions provided by them for further details. Note that the properties in Lemma 2 are not used in the existing proof of GP-UCB in [srinivas10gaussian]. As described in the next section, we analyze the realized input sequence $\mathbf{X}_{T}$ of GP-UCB by relating it to conditions in Lemma 2.
@@
We briefly summarize the existing analysis of GP-UCB (Algorithm 1) provided by srinivas10gaussian. Based on Assumptions 1 and 2, we can construct the high-probability confidence bound of the underlying function value $f(\bm{x})$ for each $\bm{x}$ and $t\in\mathbb{N}_{+}$ through the posterior distribution of $f(\bm{x})$. Specifically, by choosing a properly designed finite representative input set $\mathcal{X}_{t}\subset\mathcal{X}$ and taking into account the discretization error with Lemma 1, srinivas10gaussian showed the following events hold simultaneously with probability at least $1-\delta$:
@@
    Discretization error. The discretization error arising from $\mathcal{X}_{t}$ is at most $1/t^{2}$. Namely, $|f(\bm{x})-f([\bm{x}]_{t})|\leq 1/t^{2}$ holds for any $\bm{x}\in\mathcal{X}$ and $t\in\mathbb{N}_{+}$, where $\bm{x}]_{t}$ denotes one of the closest points of $\bm{x}$ on $\mathcal{X}_{t}$.
@@
In the above expression, the upper bound $\sum_{t=1}^{T}f(\bm{x}^{\ast})-f([\bm{x}^{\ast}]_{t})\leq\sum_{t=1}^{T}1/t^{2}\leq\pi^{2}/6$ follows from the second event (discretization error). The inequality $\sum_{t=1}^{T}f([\bm{x}^{\ast}]_{t})-f(\bm{x}_{t})\leq 2\beta_{T}^{1/2}\sum_{t=1}^{T}\sigma(\bm{x}_{t};\mathbf{X}_{t-1})$ also follows from the first event (confidence bound) and the definition of $\bm{x}_{t}$. See the proof of Theorem 2 in [srinivas10gaussian] for details. The above inequality suggests that the regret upper bound of GP-UCB depends on the sum of the posterior standard deviations $\sum_{t=1}^{T}\sigma(\bm{x}_{t};\mathbf{X}_{t-1})$. srinivas10gaussian provides the upper bound of this term by leveraging the information gain $I(\mathbf{X}_{T})$ as follows:
@@
where $C=\frac{2}{\ln(1+\sigma^{-2})}$. From Eqs. (7) and (8), we conclude that the regret upper bound of GP-UCB is $O\left(\sqrt{\beta_{T}T\gamma_{T}(\mathcal{X})}\right)$ with probability at least $1-\delta$. By combining the explicit upper bound of $\gamma_{T}(\mathcal{X})$ [srinivas10gaussian, vakili2021information], we also obtain $O\left(\sqrt{T\ln^{d+2}T}\right)$ and $\widetilde{O}\left(T^{\frac{\nu+d}{2\nu+d}}\right)$ regret upper bounds for SE and Matérn kernels, respectively.